Howard Bowen is a large-scale cotton farmer. The land and machinery he owns has a
current market value of $4 million. Bowen owes his local bank $3 million. Last year,
Bowen sold $5 million worth of cotton. His variable operating costs were $4.5 million;
accounting depreciation was $40,000, although the actual decline in value of Bowen’s
machinery was $60,000 last year. Bowen paid himself a salary of $50,000, which is not
considered part of his variable operating costs. Interest on his bank loan was $400,000.
If Bowen worked for another farmer or a local manufacturer, his annual income would be
about $30,000. Bowen can invest any funds that would be derived, if the farm sold, to
earn 10 percent annually. (Ignore taxes.)
a. Compute and interpret Bowen’s accounting profits.
b. Compute and interpret Bowen’s economic profits.
1
Expert's answer
2014-09-23T11:56:20-0400
Land and machinery - $4 million. Bowen owes his local bank $3 million. Sold $5 million worth of cotton. VC = $4.5 million, depreciation = $40,000, actual decline in value = $60,000 last year, salary = $50,000, Interest paid = $400,000. Opportunity salary = $30,000. Bowen can invest funds to earn 10 percent annually.Economic profit consists of revenue minus implicit (opportunity) and explicit (monetary) costs; accounting profit consists of revenue minus explicit costs.a. Accounting profits = 5,000,000 - 4,500,000 - 40,000 - 50,000 - 400,000 = $10,000b. Economic profit = Accounting profit - Opportunity costs = 10,000 - (60,000 - 40,000) - 30,000 - 0.1*4,000,000 = -$440,000